OpenStudy (richyw):

Find the centroid of the plane region bounded by the curves \(y=0\) and \(y=x^2-4\). Assume that the density is \(\delta =1\)

OpenStudy (richyw):

so I know that this is symmetric and therefor \(\bar{x}=0\) and since the density is zero \[m=\int^2_{-2}x^2-4\,dx=-\frac{3}{32}\]

OpenStudy (richyw):

so then I did \[\bar{y}=-\frac{3}{32}\int^2_{-2} \int^0_{x^2-4}y\,dy\,dx\]

OpenStudy (anonymous):

Then did something go wrong?

OpenStudy (richyw):

is that correct so far?

OpenStudy (anonymous):

well except for when you said "density is zero"

OpenStudy (richyw):

oops I meant one

OpenStudy (richyw):

I must have done the integration wrong then, thanks!

OpenStudy (anonymous):

well not really If you're finding M then you pretty much have everything right but the -1

OpenStudy (anonymous):

Your actual area should be \[y_1 - y_2 = 0-(x^2-4)\] Which comes out to 3/32, if I trust your integration.

OpenStudy (anonymous):

See, it's not the function that you put in to find M. You were doing \[\int\limits_{x_1}^{x_2}y_2dx\] Where it should be \[\int\limits_{x_1}^{x_2}\int\limits_{y_1}^{y_2}\delta(x, y)dydx = \int\limits_{x_1}^{x_2}\int\limits_{y_1}^{y_2}1dydx =\int\limits_{x_1}^{x_2}y_2-y_1dx \] (I messed up the subscripts in my above answer) The point is, mass should be positive.

OpenStudy (anonymous):

And, now that I check your integration, the answer is actually M = 32/3

OpenStudy (anonymous):

But you were right regarding the symmetry argument (which only holds true since \[\delta(x, y)\] is symmetric, too). Thus we know xbar = 0

OpenStudy (anonymous):

@richyw i found two minor mistakes . first the formula you wrote is for X bar nor for Y bar second integral should be multiplied by -32/3 not -3/32

OpenStudy (anonymous):

lol sami late for the party?

OpenStudy (anonymous):

yes m late..lol

OpenStudy (richyw):

alright thank you. I have the answer now!

OpenStudy (anonymous):

Anyways, your formula for ybar is right, with my additions: \[y_{centroid} = \int\limits_{x_1}^{x_2}\int\limits_{y_1}^{y_2}y \times \delta(x, y)dydx/M\] (Note that we've got ourselves a type 2 here, so we integrate with y on the inside) Substitute density for one: \[ \int\limits\limits_{x_1}^{x_2}\int\limits\limits_{y_1}^{y_2}y \times dydx/(32/3) = (3/32)\int\limits\limits_{x_1}^{x_2}\int\limits\limits_{y_1}^{y_2}y \times dydx\] FToC:\[(3/64)\int\limits\limits\limits_{x_1}^{x_2}y^2|_{x^2-4}^{0} dx = (-3/64)\int\limits\limits\limits_{x_1}^{x_2}(x^2-4)^2 dx \]

OpenStudy (anonymous):

And then it's just first semester calc from there!